# Correcting the Mythos of KL-Regularization: Direct Alignment without Overoptimization via $\chi^{2}$-Preference Optimization

### v2

@@ 6.3 Theoretical Analysis of Iterative $\boldsymbol{\chi}$PO
The second result concerns the setting where all policies are covered by $\pi_{\mathsf{ref}}$ and is easier to interpret. Indeed, if all $\pi\in\Pi$ satisfy $D_{\chi^{2}}\left(\pi\;\|\;{}\pi_{\mathsf{ref}}\right)\leq{}C^{\star}$, then $\mathsf{subopt}(\widehat{\pi},C^{\star})=0$, which implies that we can learn an $\varepsilon$-approximate minimizer using $\widetilde{O}(C^{\star}/\varepsilon^{2})$ samples. Thus, we obtain a guarantee based on unilateral concentrability (Cui and Du, 2022), which is a stronger condition, i.e., we always have $\max_{\pi}D_{\chi^{2}}\left(\pi\;\|\;{}\pi_{\mathsf{ref}}\right)\leq C_{\mathsf{uni}}$. However, per the above discussion, the first part of Theorem 6.2 is stronger than results based on unilateral concentrability and hints at a new notion of coverage for the general preference setting. Lastly, we remark that the parameter $V_{\mathsf{max}}$ only affects $\sqrt{1/m}$ in Theorem 6.2, so this dependence on this parameter can be mitigated using unlabeled data.
@@
Proof sketch
We now sketch the proof of Theorem 6.2. For a parameter $C\geq 1$, let $\widetilde{\pi}_{C}$ be the best response to $\widehat{\pi}$, restricted to the set $\Pi_{C}$:
Let $\widetilde{\pi}=\widetilde{\pi}_{\infty}$ be the global best response to $\widehat{\pi}$. Recall that $\overline{r}^{t}(x,a)\vcentcolon={}\mathbb{E}_{b\sim\pi^{t}(x)}[\widehat{\ell}(x,a,b)]$, and let $g_{x}(p)\vcentcolon=\beta D_{f_{\chi_{\textsf{mix}}}}\left(p\,\|\,\pi_{\mathsf{ref}}(x)\right)$ denote the (context-wise) mixed $\chi^{2}$-divergence. Due to the skew symmetry of $\ell^{\star}$, we only need to bound $\ell^{\star}(\widetilde{\pi},\widehat{\pi})$, which can be decomposed into the following terms:
$~{}~{}~{}~{}~{}~{}+\underbrace{\frac{1}{T}\sum_{t=1}^{T}(\overline{r}^{t}(\widetilde{\pi}_{C})-\widehat{r}^{t}(\widetilde{\pi}_{C}))}_{(3)}+\underbrace{\frac{1}{T}\sum_{t=1}^{T}(\widehat{r}^{t}(\pi^{t})-\overline{r}^{t}(\pi^{t}))}_{(4)},$
where we denote $\overline{r}^{t}(\pi):=\mathbb{E}_{x\sim\rho,a\sim\pi(x)}[\overline{r}^{t}(x,a)]$. The decomposition utilizes the fact that $\overline{r}^{t}(\pi^{t})=0$ and $\overline{r}^{t}(\widetilde{\pi}_{C})=\widehat{\ell}(\widetilde{\pi}_{C},\pi^{t})$. This implies that we only need to bound terms (1), (2), (3), and (4) to upper bound the gap of $\widehat{\pi}$. Note that term (2) is the estimation error of $\widehat{\ell}$ under the distribution $x\sim\rho,a\sim\widetilde{\pi}_{C}(x),b\sim\pi^{t}(x)$, and we can bound it by using similar techniques to the proof of Theorem 3.1. Terms (3) and (4) can be bounded by standard martingale concentration arguments. The challenge of the analysis lies in controlling term (1), because Algorithm 2 applies the $\chi$PO update (with function approximation) rather than the idealized per-state policy mirror descent update, which prevents us from utilizing existing techniques in the literature (Zhan et al., 2023a; Lan, 2023). We solve this difficulty by using the following key lemma:
Lemma 6.1.
For any $C\geq 0$, we have that for all policies $\pi\in\Pi_{C}$,
$\frac{1}{T}\sum_{t=1}^{T}\left(\widehat{r}^{t}(\pi)-\widehat{r}^{t}(\pi^{t})\right)$ $\lesssim\frac{\beta C}{\eta T}+\beta C+\frac{\eta}{\beta}-\frac{1}{T}\sum_{t=1}^{T+1}\mathbb{E}_{x\sim\rho}[g_{x}(\pi^{t})]$
$~{}~{}~{}~{}~{}~{}+\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}_{x\sim\rho}\left[\left\langle\widehat{r}^{t}(x,\cdot)-G^{t}(\pi^{t+1},x,\cdot),\pi(x)-\pi^{t+1}(x)\right\rangle\right],$
where $G^{t}(\pi,x,a):=\beta\left((1+\frac{1}{\eta})\phi\left(\frac{\pi(a|x)}{\pi_{\mathsf{ref}}(a|x)}\right)-\frac{1}{\eta}\phi\left(\frac{\pi^{t}(a|x)}{\pi_{\mathsf{ref}}(a|x)}\right)\right)$.
With this lemma, we only need to bound the quantity $\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}_{x\sim\rho}\left[\left\langle\widehat{r}^{t}(x,\cdot)-G^{t}(\pi^{t+1},x,\cdot),\pi(x)-\pi^{t+1}(x)\right\rangle\right]$ in order to bound term (1). However, this term can be upper bounded by $\mathbb{E}_{x\sim\rho,a\sim\pi(x),b\sim\pi_{\mathsf{ref}}(x)}\Big{[}\Big{|}\widehat{r}^{t}(x,a)-\widehat{r}^{t}(x,b)-f^{\beta,\eta}_{\pi^{t+1},\pi^{t}}(x,a,b)\Big{|}\Big{]}$ and $\mathbb{E}_{x\sim\rho,a\sim\pi^{t+1}(x),b\sim\pi_{\mathsf{ref}}(x)}\Big{[}\Big{|}\widehat{r}^{t}(x,a)-\widehat{r}^{t}(x,b)-f^{\beta,\eta}_{\pi^{t+1},\pi^{t}}(x,a,b)\Big{|}\Big{]}$, which can then be bounded by applying similar techniques to those in the proof of Lemma E.3. This concludes the proof.
@@ 7 Discussion
Our work gives the first practical, general-purpose algorithm for offline alignment that enjoys provable robustness to overoptimization and sample complexity guarantees based on single-policy concentrability. Conceptually, our results contribute to a growing body of research that highlights the statistical benefits of $\chi^{2}$-divergence for reinforcement learning (Wang et al., 2024; Gabbianelli et al., 2024; Amortila et al., 2024), and offer an example of fruitful interplay between reinforcement learning theory and language modeling. From this perspective, we expect that our analysis techniques and algorithm design ideas will find broader use.
We are evaluating the empirical performance of $\chi$PO in large language models as part of ongoing work. Natural technical directions raised by our paper include (i) developing a tight understanding of minimax sample complexity and instance-optimality for offline alignment with general policy classes; (ii) understanding the tightest possible problem-dependent sample complexity guarantees for offline alignment with general preference models (in light of our lower bounds in Section 6); and (iii) extending our techniques to reinforcement learning settings beyond offline alignment (e.g., general Markov decision processes).
@@ Acknowledgements
We thank Qinghua Liu for several helpful discussions.
@@ Appendix A Additional Related Work
Much of prior theoretical work on offline alignment considers algorithms that are tailored to linearly parameterized policies (Zhu et al., 2023; Zhan et al., 2023b; Li et al., 2023; Xiong et al., 2023). For general policy classes, Ye et al. (2024) provide an algorithm that achieves sample complexity guarantees based on single-policy concentrability, but the algorithm requires computation of an uncertainty bonus which cannot be implemented faithfully for large language models. Ji et al. (2024) provide an algorithm that achieves single-policy concentrability using self-play, but their approach requires the non-standard realizability assumption that for all $\pi\in\Pi$, there exists $\pi^{\prime}\in\Pi$ such that $r(x,a)=\beta\log\frac{\pi(a\mid{}x)}{\pi^{\prime}(a\mid{}x)}-Z_{\pi,\pi^{\prime}}(x)$ for some function $Z_{\pi,\pi^{\prime}}(x)$ that depends on $x$, but not the action $a$. In addition, their algorithm is iterative, and requires solving a DPO-like objective many times (roughly $1/\varepsilon^{2}$ iterations are required to achieve accuracy $\varepsilon$). Most relevant to our work, Liu et al. (2024); Cen et al. (2024); Fisch et al. (2024) propose solving the appealingly simple DPO + SFT objective in Eq. 5. As we discuss in detail in Section A.1, this objective fails to achieve single-policy concentrability unless non-standard convexity assumptions on the policy class or reward model class hold.
A number of other works consider the hybrid setting for alignment where—in addition to offline preference data from $\pi_{\mathsf{ref}}$, the algorithm has access to online feedback (Xiong et al., 2023; Gao et al., 2024; Chang et al., 2024; Song et al., 2024). While it is straightforward to achieve guarantees based on single-policy concentrability in this setting, this is a stronger feedback model than what we consider, and is not always realistic. Our work is also complementary to fully online alignment, which dispenses with coverage conditions entirely but requires active exploration (Xu et al., 2020; Novoseller et al., 2020; Pacchiano et al., 2021; Wu and Sun, 2023; Zhan et al., 2023c; Chen et al., 2022; Wang et al., 2023b; Du et al., 2024; Das et al., 2024; Ye et al., 2024; Xie et al., 2024; Cen et al., 2024).
@@
Other related works include (i) Duan et al. (2020) show that $\chi^{2}$-divergence plays a fundamental role in offline RL with linear function approximation; (ii) Zhan et al. (2022) use $\chi^{2}$-regularization to provide guarantees based on single-policy concentrability for an offline RL method based on weight function learning; and (iii) Amortila et al. (2024) provide online RL algorithms that explore by directly minimizing an exploration objective based on $\chi^{2}$-divergence. We mention in passing that a number of recent works apply $\chi^{2}$-regularization (Zhu et al., 2020; Lee et al., 2021; Ma et al., 2022a, b; Zhu and Zhang, 2024) empirically.
@@
Our work uses DPO (Rafailov et al., 2023) as a starting point. Many algorithms prior works have built upon DPO with the aim of addressing specific shortcomings, including Liu et al. (2023); Tang et al. (2024); Azar et al. (2024); Rosset et al. (2024); Chen et al. (2024); Wu et al. (2024); Tajwar et al. (2024). Closely related, there is a large body of research that attempts to understand and mitigate overoptimization in offline alignment from a purely empirical perspective (Michaud et al., 2020; Tien et al., 2022; Coste et al., 2023; Dong et al., 2023; Eisenstein et al., 2023; Gao et al., 2023; Moskovitz et al., 2023; Pal et al., 2024; Rita et al., 2024; Rafailov et al., 2024a; Zhang et al., 2024).
@@ A.1 Detailed Comparison to DPO + SFT
In this section, we give additional background on the suboptimality of the DPO + SFT objective in Eq. 5. Let $\beta>0$ be the KL-regularization parameter and $\alpha>0$ be an optimism parameter. Consider the setting in which $\Pi=\left\{\pi_{r}(a\mid{}x)=\pi_{\mathsf{ref}}(a\mid{}x)\exp(\beta^{-1}(r(x,a)-Z_{r}(x)))\mid{}r\in\mathcal{R}\right\}$ for a given reward function class $\mathcal{R}\subset(\mathcal{X}\times\mathcal{A}\to\mathbb{R})$. Liu et al. (2024); Cen et al. (2024); Fisch et al. (2024) propose solving (variants of) the objective
@@
highlighting the fact that $\chi$PO meaningfully adapts to single-policy concentrability even when the technical conditions required by DPO+SFT do not hold; see also Section 4. We find this conclusion to be somewhat surprising, as Xie et al. (2024) show that an optimistic counterpart to Eq. 31, which negates the SFT term, enjoys strong guarantees for online alignment with general policy classes without requiring convexity conditions.
@@
Appendix B Applying $\boldsymbol{\chi}$PO to the Token-Level MDP

### v3

@@ 7.3 Theoretical Analysis of Iterative $\boldsymbol{\chi}$PO
The second result concerns the setting where all policies are covered by $\pi_{\mathsf{ref}}$ and is easier to interpret. Indeed, if all $\pi\in\Pi$ satisfy $D_{\chi^{2}}\left(\pi\;\|\;{}\pi_{\mathsf{ref}}\right)\leq{}C^{\star}$, then $\mathsf{subopt}(\widehat{\pi},C^{\star})=0$, which implies that we can learn an $\varepsilon$-approximate minimizer using $\widetilde{O}(C^{\star}/\varepsilon^{2})$ samples. Thus, we obtain a guarantee based on unilateral concentrability (Cui and Du, 2022), which is a stronger condition, i.e., we always have $\max_{\pi}D_{\chi^{2}}\left(\pi\;\|\;{}\pi_{\mathsf{ref}}\right)\leq C_{\mathsf{uni}}$. However, per the above discussion, the first part of Theorem 7.2 is stronger than results based on unilateral concentrability and hints at a new notion of coverage for general preferences. Lastly, we remark that the parameter $V_{\mathsf{max}}$ only affects $\sqrt{1/m}$ term in Theorem 7.2, so dependence on this parameter can be mitigated using unlabeled data.
@@ 8 Discussion
Our work gives the first practical, general-purpose algorithm for offline alignment with provable robustness to overoptimization and sample complexity guarantees based on single-policy concentrability. Conceptually, our results contribute to a growing body of research that highlights the statistical benefits of $\chi^{2}$-divergence for reinforcement learning (Wang et al., 2024; Gabbianelli et al., 2024; Amortila et al., 2024), and offer an example of fruitful interplay between reinforcement learning theory and language modeling. From this perspective, we expect that our analysis techniques and algorithm design ideas will find broader use.
Natural technical directions raised by our paper include (i) developing a tight understanding of minimax sample complexity and instance-optimality for offline alignment with general policy classes; (ii) understanding the tightest possible problem-dependent sample complexity guarantees for offline alignment with general preference models (in light of lower bounds in Section 7); and (iii) extending our techniques to reinforcement learning settings beyond offline alignment (e.g., general MDPs). We look forward to studying these questions in future work.
@@ Acknowledgements
We thank Qinghua Liu, Zhaolin Gao, and Yuda Song for several helpful discussions. WS acknowledges funding support from NSF IIS-2154711, NSF CAREER 2339395, DARPA LANCER: LeArning Network CybERagents.
@@ Appendix A Additional Related Work
Much of prior theoretical work on offline alignment considers algorithms that are tailored to linearly parameterized policies (Zhu et al., 2023; Li et al., 2023; Xiong et al., 2023), while others are not efficiently implementable, e.g., as they require solving min-max problems over a version space (Zhan et al., 2023a). For general policy classes, Ye et al. (2024) provide an algorithm that achieves sample complexity guarantees based on single-policy concentrability, but the algorithm requires computation of an uncertainty bonus which cannot be implemented faithfully for large language models. Ji et al. (2024) provide an algorithm that achieves single-policy concentrability using self-play, but their approach requires the non-standard realizability assumption that for all $\pi\in\Pi$, there exists $\pi^{\prime}\in\Pi$ such that $r(x,a)=\beta\log\frac{\pi(a\mid{}x)}{\pi^{\prime}(a\mid{}x)}-Z_{\pi,\pi^{\prime}}(x)$ for some function $Z_{\pi,\pi^{\prime}}(x)$ that depends on $x$, but not the action $a$. In addition, their algorithm is iterative, and requires solving a DPO-like objective many times (roughly $1/\varepsilon^{2}$ iterations are required to achieve accuracy $\varepsilon$). Most relevant to our work, Liu et al. (2024); Cen et al. (2024); Fisch et al. (2024) propose solving the appealingly simple DPO + SFT objective in Eq. 5. As we discuss in detail in Section A.1, this objective fails to achieve single-policy concentrability unless non-standard convexity assumptions on the policy class or reward model class hold.
A number of other works consider the hybrid setting for alignment where—in addition to offline preference data from $\pi_{\mathsf{ref}}$, the algorithm has access to online feedback (Xiong et al., 2023; Gao et al., 2024; Chang et al., 2024; Song et al., 2024). While it is straightforward to achieve guarantees based on single-policy concentrability in this setting, this is a stronger feedback model than what we consider, and is not always realistic. Our work is also complementary to fully online alignment, which dispenses with coverage conditions entirely but requires active exploration (Xu et al., 2020; Novoseller et al., 2020; Pacchiano et al., 2021; Wu and Sun, 2023; Zhan et al., 2023b; Chen et al., 2022; Wang et al., 2023b; Du et al., 2024; Das et al., 2024; Ye et al., 2024; Xie et al., 2024; Cen et al., 2024).
@@
Other related works include (i) Duan et al. (2020) show that $\chi^{2}$-divergence plays a fundamental role in offline RL with linear function approximation; (ii) Zhan et al. (2022) use $\chi^{2}$-regularization to provide guarantees based on single-policy concentrability for an offline RL method based on weight function learning; and (iii) Amortila et al. (2024) provide online RL algorithms that explore by directly minimizing an exploration objective based on $\chi^{2}$-divergence. We mention in passing that a number of recent empirical works apply $\chi^{2}$-regularization (Zhu et al., 2020; Lee et al., 2021; Ma et al., 2022a, b; Zhu and Zhang, 2024) to reinforcement learning in embodied domains. Lastly, Cesa-Bianchi et al. (2017) prove lower bounds against the softmax policy distribution, but in the context of online exploration for online RL. While this is different problem setting than ours, their construction may be in similar in spirit to our lower bound against KL-regularization in offline reinforcement learning (Proposition A.1).
@@
Our work uses DPO (Rafailov et al., 2023) as a starting point. Many prior works have built upon DPO with the aim of addressing specific shortcomings, including Liu et al. (2023); Tang et al. (2024); Azar et al. (2024); Rosset et al. (2024); Chen et al. (2024); Wu et al. (2024); Tajwar et al. (2024). Closely related, there is a large body of research that attempts to understand and mitigate overoptimization in offline alignment from a purely empirical perspective (Michaud et al., 2020; Tien et al., 2022; Coste et al., 2023; Dong et al., 2023; Eisenstein et al., 2023; Gao et al., 2023; Moskovitz et al., 2023; Pal et al., 2024; Rita et al., 2024; Rafailov et al., 2024a; Zhang et al., 2024).
@@ A.1 Detailed Comparison to DPO + SFT
In this section, we give additional background on the suboptimality of the DPO + SFT objective in Eq. 5. Let $\beta>0$ be the KL-regularization parameter and $\alpha>0$ be an optimism parameter. Consider the setting in which $\Pi=\left\{\pi_{r}(a\mid{}x)=\pi_{\mathsf{ref}}(a\mid{}x)\exp(\beta^{-1}(r(x,a)-Z_{r}(x)))\mid{}r\in\mathcal{R}\right\}$ for a reward class $\mathcal{R}\subset(\mathcal{X}\times\mathcal{A}\to\mathbb{R})$. Liu et al. (2024); Cen et al. (2024); Fisch et al. (2024) propose solving (variants of) the objective
@@
highlighting the fact that $\chi$PO meaningfully adapts to single-policy concentrability even when the technical conditions required by DPO+SFT do not hold; see also Section 4. We find this conclusion to be somewhat surprising, as Xie et al. (2024) show that an optimistic counterpart to Eq. 31, which negates the SFT term, enjoys strong guarantees for online alignment with general policy classes without requiring convexity.
@@
Appendix B Sample Complexity Guarantees for $\boldsymbol{\chi^{2}}$-RLHF
Algorithm
Algorithm 3 $\chi^{2}$-RLHF
Theoretical guarantees
Assumption B.1.
Assumption B.2.
Below is our main sample complexity guarantee for $\chi^{2}$-RLHF. While it is stated for a fixed, $\beta$-dependent smoothing parameter for compactness, the general version of this result (Theorem I.1) allows for general $\eta$.
Theorem B.1.
Let $\beta>0$ be given, and suppose Assumptions B.1 and B.2 hold any $\eta\in\left[0,\frac{\beta}{8R_{\mathsf{max}}}\right]$. With probability at least $1-\delta$, $\chi^{2}$-RLHF (Algorithm 3) produces a policy $\widehat{\pi}$ such that for all policies $\pi^{\star}$ simultaneously, we have
$J(\pi^{\star})-J(\widehat{\pi})$
$\lesssim R_{\mathsf{max}}e^{2R_{\mathsf{max}}}\cdot\sqrt{\frac{\mathcal{C}^{\pi^{\star}}\log(|\mathcal{R}|/\delta)}{n}}+\beta\cdot\mathcal{C}^{\pi^{\star}}+\beta^{-1}\cdot\frac{R_{\mathsf{max}}^{2}e^{4R_{\mathsf{max}}}\log(|\mathcal{R}|/\delta)}{n}+R_{\mathsf{max}}\sqrt{\frac{\log(|\Pi|/\delta)}{n_{\mathsf{x}}}}+\varepsilon_{\mathrm{opt}}.$
Above, we see that $\chi^{2}$-RLHF, like $\chi$PO, has sample complexity that scales only with the single-policy concentrability coefficient $\mathcal{C}^{\pi^{\star}}$, and holds for all comparator policies $\pi^{\star}$ simultaneously. Since the choice of $\beta$ induces a similar bias-overoptimization tradeoff in the first statement of Theorem B.1 as it did in Theorem 3.1 for $\chi$PO, we focus our discussion on the guarantee for a tuned choice of $\beta$ (Eq. 34). The first term in Eq. 34 accounts for the reward estimation error (2) and scales with $\mathcal{C}^{\pi^{\star}}$; as before, this accounts for how well rewards estimated from $\pi_{\mathsf{ref}}$ transfer to other candidate policies. The second term in Eq. 34 accounts for the statistical error from sampled contexts used in 4 for policy optimization. In particular, it is possible to drive this term to be much smaller than the first by using a larger unlabeled context dataset, which is typically far cheaper to acquire.
Computationally efficiency
Theorem B.1 bounds the sample complexity of $\chi^{2}$-RLHF under the assumption that we can solve 4 up to $\varepsilon_{\mathrm{opt}}$-accuracy. This is a purely computational problem, and in practice it can be solved using policy gradient methods such as PPO.
Comparison to $\boldsymbol{\chi}$PO
Unlike $\chi$PO (Theorem 3.1), Theorem B.1 has no dependence on the parameter $V_{\mathsf{max}}$ or quantities such as $\frac{\pi}{\pi_{\mathsf{ref}}}\leq\max_{\pi}\mathcal{C}_{\infty}^{\pi}$. We primarily attribute this to the fact that $\chi^{2}$-RLHF uses an explicit reward function class $\mathcal{R}$, and normalizing or clipping it to the reward range $R_{\mathsf{max}}$ is both natural and routinely done in practice (Shah et al., 2015; Christiano et al., 2017; Ouyang et al., 2022). In comparison, the implicit reward models induced by the policy class $\Pi$ in $\chi$PO can have larger range, and clipping the policy class in $\chi$PO directly, e.g., so that $|\beta\phi(\frac{\pi}{\pi_{\mathsf{ref}}})|$ is bounded, is misguided, because the policy class may lose realizability (Assumption 3.1). This is because $r^{\star}(x,a)=\beta\phi\Big{(}\frac{\pi^{\star}_{{\scriptscriptstyle\beta}}(a|x)}{\pi_{\mathsf{ref}}(a|x)}\Big{)}+Z_{\beta,r^{\star}}(x)$, and the normalization factor $Z_{\beta,r^{\star}}$ cannot be reasonably accounted for when clipping $\Pi$. While the $V_{\mathsf{max}}$ (Assumption 3.2) parameter involves pairs of action probabilities, and thereby sidesteps the normalization constant issue, it may not always be practical to modify $\Pi$ so that $V_{\mathsf{max}}$ is bounded, since this would require checking all pairs of each policy’s action probabilities.
However, using an explicit reward function class alone is not enough. As discussed previously, when we move from implicit to explicit $\chi^{2}$-regularization, incorporating the smoothing parameter $\eta$ in Eq. 32 is essential to avoid statistical errors due to policies with large density ratios when we approximate the $\chi^{2}$-regularizer with empirical data. A careful choice of $\eta=\beta/R_{\mathsf{max}}$ in Theorem B.1 balances the benefits of clipping against the bias it introduces. Without smoothing (i.e., $\eta=0$), a guarantee that depends on $\max_{\pi}\mathcal{C}_{\infty}^{\pi}$ for $\chi^{2}$-RLHF would be unavoidable, since the sample complexity must scale with the range of the problem, which grows with the magnitude of the regularizer. See Corollary I.2 in Appendix I for a guarantee in the case where $\eta=0$, which highlights this.
Appendix C Experiment details
Dataset and models
For training, we use trl-internal-testing/tldr-preference-trl-style, with 92.9K train samples and 83.8K validation samples. The reference policy $\pi_{\mathsf{ref}}$ is the Pythia-1b model (Biderman et al., 2023) pre-trained on SFT data (cleanrl/EleutherAI_pythia-1b-deduped__sft__tldr from Huang et al. (2022)), and performance is measured via winrate against a baseline, as judged by GPT-4o. All parameters that are not algorithm-specific, such as the learning rate, are shared by both $\chi$PO and DPO in order to ensure a fair comparison.
Training details
Our implementation of $\chi$PO is built upon the DPO trainer from Transformer Reinforcement Learning (TRL) (von Werra et al., 2020). $\chi$PO comes with strong robustness and theoretical properties, but the policy ratios can sometimes introduce instability in training. In practice, we have observed that better stability and performance can be achieved by utilizing the (more general form) link function $\widetilde{\phi}(z)\vcentcolon={}\exp\left(\mathsf{clip}_{[-88,20]}\left(\alpha\cdot\log z\right)\right)+\gamma\cdot\log z$ in Algorithm 1, and performing a small grid search over additional parameters $\alpha=\{\frac{1}{4},1\}$ and $\gamma=\{0.1,1\}$ for a fixed $\beta$.
We briefly discuss each parameter in turn. The mixing parameter $\gamma$ controls the relative ratios of KL- and $\chi^{2}$-regularization, our analysis in Section F.1 shows that Theorem 3.1 holds more generally for $\gamma\in(0,1]$ (see Theorem F.1). Next, ignoring clipping, $\alpha\in(0,1]$ in $\widetilde{\phi}$ implements regularization with the $(1+\alpha)$-divergence (or Renyi divergence), which is an $f$-divergence that is stronger than KL-regularization but weaker than $\chi^{2}$-regularization (Van Erven and Harremos, 2014), and also carries single-policy concentrability guarantees (although with a slower-rate dependence on sample size $n$). For example, $\alpha=\frac{1}{4}$ corresponds to the link function $\phi(z)=(z)^{1/4}+\gamma\log z$, which is easier to optimize than the link function $\phi(z)=z+\gamma\log z$ (corresponding to $\alpha=1$) induced by $\chi^{2}$-regularization, given the potentially large magnitude of $z=\frac{\pi}{\pi_{\mathsf{ref}}}$. Though we do not write out the analysis here, the methods used to prove the sample complexity of $\chi$PO (Theorem 3.1) can be used to prove analogous guarantees for regularization with $\alpha$-divergences, which will have slightly worse statistical rates.
Lastly, we provide some additional explanation for the clipping operation. We observed that torch.exp is prone to underflow when $\log\frac{\pi}{\pi_{\mathsf{ref}}}$ is very negative, and clipping the upper range to 20 can help reduce numerical instabilities. Clipping in such a manner is supported by our analysis in Proposition 4.2, which shows that $\frac{\pi^{\star}}{\pi_{\mathsf{ref}}}\leq 1+\frac{R_{\mathsf{max}}}{\beta}$ (though technically we do not know $R_{\mathsf{max}}$). The parameters for all experiments are displayed in Table 2.
Table 2: Parameter settings in TL;DR summarizion
Algorithm Parameters
DPO batch size: 64
learning rate: 1e-6
scheduler: cosine
optimizer: adamw
$\chi$PO batch size: 64
clip range: [-88, 20]
learning rate: 1e-6
scheduler: cosine
optimizer: adamw
$\beta=0.05$, 1 epoch $\alpha:1.25,\gamma:1.0$
$\beta=0.05$, 2 epochs $\alpha:2.00,\gamma:1.0$
$\beta=0.05$, 4 epochs $\alpha:1.25,\gamma:0.1$
$\beta=0.005$, all epochs $\alpha:1.25,\gamma:0.1$
Generation details
For winrate evaluation, we use greedy, temperature 0, decoding. For computation of the KL divergence, we sample from the model with temperature 1. The maximum prompt length is 512, and the maximum response length is 200. We use the standard generation prompt “TL;DR:” (Gao et al., 2024).
Evaluation of performance
The performance of each algorithm is measured via winrate against responses in the SFT dataset, as measured by GPT-4o (global standard). The winrate is computed on a subset of 512 prompts from the SFT validation set (trl-internal-testing/tldr-preference-sft-trl-style), and the order of the model and reference responses are randomized each round.
Appendix D Applying $\boldsymbol{\chi}$PO to the Token-Level MDP